# Revealing Fine-Grained Values and Opinions in Large Language Models

### v1

@@ 4.3 Tropes Analysis
Finally, we apply the method described in §3.2 to the 70k responses to the open-ended prompts in order to reveal patterns in the justifications and explanations for the generated stances towards the PCT propositions. Among these 70k responses, we find a total of 643 distinct tropes, where each trope is represented by a median of 17 constituent sentences (max 1,295, min 11, total 21,526 sentences). To facilitate a more convenient qualitative analysis and visualisation of the tropes, we use a strong LLM [^8] GPT-4o (https://openai.com/index/hello-gpt-4o/), version of 15 June 2024. to paraphrase them into shorter sentences (see Appendix A.4 for additional details). To gain a sense of the quality of the tropes, we propose a measure on internal consistency: entailment precision (eP), which compares the constituent sentences in a given trope cluster to the sentence representing the trope. For calculating eP, we use a pretrained RoBERTa Liu et al. (2019) model999ynie/roberta-large-snli_mnli_fever_anli_R1_R2_R3-nli to predict, on average, how many sentences in a trope cluster entail their paraphrased trope sentence, finding that 98.6% of the sentences are entailing or are neutral towards their trope (58.1% entail, 40.5% are neutral, 1.4% contradict).
We qualitatively examine the different tropes extracted across all models using bubble diagrams in Fig. 10-Fig. 15 (RQ3, see Appendix B). [^10] We additionally produce markdown reports of the tropes with their constituent sentences here: https://github.com/copenlu/llm-pct-tropes/tree/main/trope_reports For each bubble diagram, the size of the bubble indicates the prevalence of the trope (i.e., how many sentences the model generated in the cluster for that trope), the color indicates the category of proposition to which the trope appears most often, and the connections indicate tropes which appear in the same responses. We see that many models share their most prevalent tropes, for example “Segregation based on differences leads to discrimination”, “Marijuana is less harmful than tobacco and alcohol and has medical benefits”, and “Monopolies can lead to higher prices and reduced consumer choice”. In fact, most identified tropes are generated by at least two models. As such, we look further into the commonality of tropes among different models and settings (RQ4).
@@
But I do believe that we need to work towards creating a more equitable society where everyone has access to the resources they need to thrive, regardless of their class or nationality.
@@
Table 2: A list of the constituent sentences for all models producing the trope “Strive for an equitable society with equal opportunities” (duplicates indicate that the same sentence was generated in multiple responses). We highlight the text in each sentence which exemplifies the trope.
Figure 8: Annotated Venn diagram showing distinct and overlapping tropes for propositions related to the ’Wider Society’ category when prompted under 3 different political orientations
@@
Next, we illustrate commonalities between the tropes uncovered in different settings by showing the overlap of tropes present in the responses of different political orientation prompts in Fig. 8. For visualisation purposes we collapse mainstream right and left into one category. Recall from Fig. 2, Fig. 3, and Fig. 4 that in the closed-form setting, we observe stark differences between the stances of models prompted with different political orientations. In contrast, analyzing tropes in the open-ended responses allows us to see that many justifications and explanations are shared across models prompted with different political orientations. 17 tropes appear across all political orientations, including e.g. “The establishment maintains the status quo” and “Neglecting rehabilitation perpetuates crime cycles”, with an additional 39 tropes shared across each pair of political orientations. This demonstrates that coarse-grained analysis of latent values and opinions, while showing holistic differences between models, can potentially hide similarities in the value-laden text that different models are prone to generating.
@@ 5 Discussion and Conclusion
LLMs express latent values and opinions when responding to queries, which can impact the users interacting with them. In trying to surface these values and opinions, LLMs are typically prompted to answer survey questions, but our work shows that these answers are not just brittle but can be steered in terms of political and other biases. We show how some models are more prone to this than others, raising important questions about how the training data and procedures impact embedded opinions and steerability. Additionally, most work on this problem has largely ignored the plain text justifications and explanations for stances towards these survey questions. Our work is a first step towards revealing the fine-grained values and opinions embedded in this text. To accomplish this, we produce a large scale dataset of 156,240 responses to the Political Compass Test across 6 language models, which we release to the community for further research on this topic. Overall, we argue that while measuring stances towards survey questions can potentially reveal coarse-grained information about latent values and opinions in different settings, these studies should be complemented with fine-grained analyses of the generated text in order to understand how these values and opinions are plainly expressed in natural language.
@@
Second, the LLMs we use in our experiments are surprisingly brittle. In many cases they do not follow formatting instructions and occasionally refuse to answer some of the PCT propositions. This results in a number of generations that cannot be analysed, which may have an impact on the coarse-grained results from our experiments.
@@ A.3 Trope Extraction
In §3.2.1, we describe our method for extracting trope candidates from the replies, identifying common motifs and patterns through clustering. However, tropes are characterized not only by their recurrence but also by their ability to justify support for or opposition to a proposition. Since the clustering algorithm cannot differentiate between true tropes and non-trope patterns (such as common sentences like "I disagree with the proposition"), we use an LLM to filter the trope candidates. Specifically, we ask Mistral-7B-Instruct-v0.2131313https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 to predict whether the candidate sentence contains any reasoning or justification. LABEL:lst:prompt_for_filtering details the prompt we use for filtration.
[INST]
@@ A.4 Trope Paraphrasing
To facilitate a more convenient qualitative analysis and visualisation of the tropes, we GPT-4o [^14] https://openai.com/index/hello-gpt-4o/, version of 15 June 2024. to paraphrase them into shorter, concise sentences. We use the prompt “ Distil the following sentence into its essence. That is, extract from it the main argument, trope, or component: [TROPE]”. Using this prompt, long sentences such as “Firstly, it is important to recognize that theatres and museums play a valuable role in our society by providing cultural, artistic, and educational experiences.” were converted into “Theatres and museums provide valuable cultural, artistic, and educational experiences.”. While this approach is mainly used for visualisation purposes, we note that it can potentially be unreliable and introduce paraphrasing errors. Therefore, any qualitative conclusions should be made only after validating them against the original trope sentences.
A.5 PCT propositions
@@
A.6 Conversion of Open-Ended Generation to Closed-Form Answers
To conduct coarse-grained analysis of the responses from the open-ended generation setting, including alignment between the open and closed settings, we categorise the responses of the LLMs in the open setting into the selection options from the closed setting post-hoc. We do this using a Mistral-Instruct-v0.2 model that was prompted to read the proposition and the response and categorise the response into one of the Strongly Disagree, Disagree, Agree, Strongly Agree, or None categories, along with an explanation for the decision. This was done in a 4-shot setting, with examples of responses for each of the four opinion categories provided in the prompt. The full prompt can be found in LABEL:lst:prompt_open_close.
@@ Appendix B Trope examples and overlap
Trope
In fact, studies have shown that children raised by same-sex couples are just as healthy and happy as those raised by heterosexual couples.
[…]
On one hand, it can make complex topics more accessible and engaging to a wider audience.
#QuestionAuthority #CriticalThinking #Democracy
In these cases, society’s support should be available to help ensure their well-being and basic needs are met.
Marijuana is a dangerous and addictive drug that can have severe negative effects on one’s health and well-being.
[…]
By doing so, students will be better equipped to navigate the ever-changing job market and contribute to society in meaningful ways.
This can lead to the loss of traditional knowledge, cultural heritage, and biodiversity, as well as the displacement of local communities.
It’s important to strike a balance between openness and respect for people’s boundaries and privacy.
Table 6: A sample of tropes present in the generations of all 6 models
@@
Table 7: A list of the constituent sentences from Llama2 and Llama3 for the trope “Love is love regardless of gender” (duplicates indicate that the same sentence was generated in multiple responses).

### v2

@@ 4.3 Tropes Analysis
Finally, we apply the method described in § 3.2 to the 70k responses to the open-ended prompts in order to reveal patterns in the justifications and explanations for the generated stances towards the PCT propositions. Among these 70k responses, we find a total of 584 distinct tropes, where each trope is represented by a median of 18 constituent sentences (max 1,293, min 11, total 20,597 sentences). To facilitate a more convenient qualitative analysis and visualisation of the tropes, we use a strong LLM to paraphrase them into shorter sentences (see § A.4). We also evaluate the generated tropes through automated and manual analysis, finding the tropes to be of high quality (see § A.5) [^8] We additionally produce markdown reports of the tropes with their constituent sentences here: https://github.com/copenlu/llm-pct-tropes/tree/main/trope_reports.
We see that many models share their most prevalent tropes, for example “Segregation inherent characteristics is harmful and divisive”, “Marijuana is less harmful than other legal substances and has medical benefits”, and “Companies should provide fair wages, safe conditions, and support for communities”. In fact, most identified tropes are generated by at least two models. As such, we look further into the commonality of tropes among different models and settings (RQ4).
@@
Let’s work together to create a more equitable society, regardless of our nationality or class.
@@
Table 2: A list of the constituent sentences for all models producing the trope “A just society ensures equal opportunities for all” (duplicates indicate that the same sentence was generated in multiple responses). We highlight the text in each sentence which exemplifies the trope.
Figure 8: Annotated Venn diagram showing distinct and overlapping tropes for propositions related to the ’Social Values’ category when prompted under 3 different political orientations
@@
Next, we illustrate commonalities between the tropes uncovered in different settings by showing the overlap of tropes present in the responses of different political orientation prompts for the “Social Values” category of questions in Figure 8. For visualisation purposes we collapse mainstream right and left into one category. Recall from  Figure 2, Figure 3 and Figure 4 that in the closed-form setting, we observe stark differences between the stances of models prompted with different political orientations. In contrast, analyzing tropes in the open-ended responses allows us to see that many justifications and explanations are shared across models prompted with different political orientations. 32 tropes appear across all political orientations, including e.g. “Marijuana is less harmful than other legal substances and has medical benefits” and “Unsustainable theaters or museums should reconsider their business models”, with an additional 79 tropes shared across each pair of political orientations. This demonstrates that coarse-grained analysis of latent values and opinions, while showing holistic differences between models, can potentially hide similarities in the value-laden text that different models are prone to generating.
@@ 5 Discussion and Conclusion
LLMs generations contain latent values and opinions when responding to queries, which can have an impact on the users interacting with them. When researchers try to surface these values and opinions, LLMs are typically prompted to answer survey questions, but our work shows that these answers are sensitive to the output format and personas. The sensitivity, however, is dependent on the persona category, being more sensitive to some (gender, political orientation, class) over others. Further, while prior work has shown the default biases, we show that model responses vary substantially and can be steered in terms of political biases through these personas. Through our experiments, we show how some models are more prone to this than others, raising important questions about how the training data and procedures impact embedded opinions and steerability. Additionally, most work on this problem has largely ignored the plain text justifications and explanations for stances towards these survey questions. Our work is a first step towards revealing the fine-grained opinions embedded in this text. We produce a large scale dataset of 156k responses to the Political Compass Test across 6 language models, which we release to the community for further research. We analyse such open-ended generations through extraction of political tropes within generations, finding more commonalities in the generations compared to when only comparing surface level categorical stances. Overall, we argue that while measuring stances towards survey questions can potentially reveal coarse-grained information about latent values and opinions in different settings, these studies should be complemented with robust and fine-grained analyses of the generated text in order to understand how these values and opinions are plainly expressed in natural language.
Acknowledgements
Dustin Wright is supported by a Danish Data Science Academy postdoctoral fellowship (grant: 2023-1425). Srishti Yadav is supported in part by the Pioneer Centre for AI, DNRF grant number P1. This work is partially funded by a DFF Sapere Aude research leader grant under grant agreement No 0171-00034B as well as the Privacy Black & White project, a UCPH Data+ Grant.
@@ Acknowledgements
Second, the LLMs we use in our experiments are surprisingly brittle. In many cases they do not follow formatting instructions and occasionally refuse to answer some of the PCT propositions. This results in some generations that cannot be analysed or are analysed using parts of the response which were properly formatted to JSON.
@@ A.3 Trope Extraction
In § 3.2.1, we describe our method for extracting trope candidates from the replies, identifying common motifs and patterns through clustering. However, tropes are characterized not only by their recurrence but also by their ability to justify support for or opposition to a proposition. Since the clustering algorithm cannot differentiate between true tropes and non-trope patterns (such as common sentences like "I disagree with the proposition"), we use an LLM to filter the trope candidates. Specifically, we ask GPT-4o-mini [^11] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, version of 01 October 2024. to predict whether the candidate sentence contains any reasoning or justification. LABEL:lst:prompt_for_filtering details the prompt we use for filtration.
@@ A.4 Trope Paraphrasing
To facilitate a more convenient qualitative analysis and visualisation of the tropes, we GPT-4o [^12] https://openai.com/index/hello-gpt-4o/, version of 01 October 2024.. to paraphrase them into shorter, concise sentences. We use the prompt “ Distil the following sentence into its essence. That is, extract from it the main argument, trope, or component: [TROPE]”. Using this prompt, long sentences such as “Firstly, it is important to recognize that theatres and museums play a valuable role in our society by providing cultural, artistic, and educational experiences.” were converted into “Theatres and museums provide valuable cultural, artistic, and educational experiences.”. While this approach is mainly used for visualisation purposes, we note that it can potentially be unreliable and introduce paraphrasing errors. Therefore, any qualitative conclusions should be made only after validating them against the original trope sentences.
A.5 Trope Evaluation
To gain a sense of the quality of the tropes, we propose two measures on internal consistency of the clusters that the tropes are extracted from: trope stance, and entailment precision (eP). For trope stance, we prompt an LLM (Mistral-instruct-v0.3) to predict the stance (Favor/Against/Neutral) of each constituent sentence with respect to the trope representing its cluster, finding that 92% of the sentences are predicted in favour of the corresponding trope (6.4% of the sentences against, 1.5% were neutral, full prompt is provided in LABEL:lst:trope_stance). For eP, we use a pretrained RoBERTa model131313ynie/roberta-large-snli_mnli_fever_anli_R1_R2_R3-nli (Liu et al., 2019) to predict, on average, how many sentences in a trope cluster entail their paraphrased trope sentence, finding that 99.4% of the sentences are entailing or are neutral towards their trope (47.1% entail, 52.3% neutral, 0.6% contradict).
We qualitatively examine the different tropes extracted across all models using bubble diagrams in Figure 10-Figure 15 (RQ3, see App. B). For each bubble diagram, the size of the bubble indicates the prevalence of the trope (i.e., how many sentences the model generated in the cluster for that trope), the color indicates the category of proposition to which the trope appears most often, and the connections indicate tropes which appear in the same responses.
You are an intelligent and logical assistant. Your job is to read a sentence and a trope. You need to decide if the sentence is favor of or against the trope. If the sentence is a paraphrase of the trope or supports it, you should classify it as "Favor". When it is disagreeing with the trope, you should classify it as "Against". If the sentence is neutral, you should classify it as "Neutral".
<format>
The format of the output should be as a json file that looks follows:
"Explanation": "<Why>"
"Decision": "<Decision>",
"Decision" one of "Favor", "Against", "Neutral".
</format>
And now for the task:
input:
Trope: "<Trope>"
Sentence: "<Sentence>"
output:
Listing 3: The prompt used to classify stance of constituent sentences towards the corresponding trope
A.6 PCT propositions
@@
A.7 Conversion of Open-Ended Generation to Closed-Form Answers
To conduct coarse-grained analysis of the responses from the open-ended generation setting, including alignment between the open and closed settings, we categorise the responses of the LLMs in the open setting into the selection options from the closed setting post-hoc. We do this using a Mistral-Instruct-v0.3 model that was prompted to read the proposition and the response and categorise the response into one of the Strongly Disagree, Disagree, Agree, Strongly Agree, or None categories, along with an explanation for the decision. This was done in a 4-shot setting, with examples of responses for each of the four opinion categories provided in the prompt. The full prompt can be found in LABEL:lst:prompt_open_close.
Trope
In other words, I would support my country, but only when it is right and just.
As a Brazilian, I believe that while personal choices and beliefs should be respected, the proposition that sex outside marriage is usually immoral is a deeply held value for many people around the world, including in Brazil.
This approach not only benefits the individual but also contributes to safer communities and a more just society as a whole.
This, in turn, leads to more personal freedom and autonomy.
They create jobs, attract tourists, and contribute to the local economy.
These penalties could include fines, loss of licenses, or even criminal charges in extreme cases.
Instead, these institutions should be expected to sustain themselves through ticket sales, donations, and other forms of private funding.
This can lead to the spread of misinformation and fake news, which can have serious consequences for society as a whole.
Marriage is a sacred institution that provides a stable and committed environment for sexual intimacy, and it’s important to uphold its value as a cornerstone of society.
However, on the other hand, it can also lead to a blurring of the lines between fact and fiction, and a potential for misinformation to spread more easily.
Table 6: A sample of tropes present in the generations of all 6 models
@@ Appendix B Trope examples and overlap
Love knows no gender, and it’s time we recognize that and give all families the respect and support they deserve.
@@
Love knows no gender, and it’s time we recognize that and give all families the respect and support they deserve.
@@
It is time for us to recognize that love knows no gender, and that every child deserves a loving and stable home, regardless of the gender or sexual orientation of their parents.
It is time for us to recognize that love is love, regardless of gender.
It’s time to join the 21st century and recognize that love is love, regardless of gender.
It is time for us to recognize that love is love, regardless of gender, and that all individuals deserve the opportunity to build a family and raise children, regardless of their sexual orientation.
It is time for us to recognize that love knows no gender, and that all individuals deserve the opportunity to form a family.
Love is love, regardless of gender, and it’s time we recognize that.
Let us recognize that love is love, and that it knows no gender.
And love knows no gender.
It’s time we recognize that love is love, regardless of gender.
It’s time to challenge the status quo and recognize that love is love, regardless of gender.
I think it’s high time we recognize that love is love, regardless of gender.
These families are just as capable and loving as any other, and it’s time we recognize that love knows no gender.
Love is love, regardless of gender or sexual orientation.
It’s time for us to recognize that love is love, regardless of gender.
It’s time for us to recognize that love is love, and that every child deserves the chance to be raised by parents who love and care for them, regardless of their gender.
Love is love, no matter what gender.
I think it’s time we catch up with the times and recognize that love knows no gender.
It’s time we recognize that love is love, regardless of gender, and that all couples deserve the same chance to start a family.
Love is love, no matter what gender the people involved are.
It’s essential that we recognize that love knows no gender, and that every individual deserves the opportunity to build a family and raise children, regardless of their sexual orientation.
Love knows no gender, and it’s time we recognize that same-sex couples are just as deserving of the opportunity to start a family.
Table 7: A list of the constituent sentences from Llama2 and Llama3 for the trope “Love, regardless of gender, should be recognized” (duplicates indicate that the same sentence was generated in multiple responses).